I am stuck on this algebra. Any help would be appreciated...
Graph f(x) = x^2+2x-6
What is the vertex? Coordinate pair.
What is the line of symmetry? x=
Is it max or min?
I don't have this chapter in my book I am using.

Answer by RAY100(1637)

y=x^2 +2x -6
.
(y+6) = x^2 +2x,,,,,,looking for form (y-k) = (x-h)^2,,,,with vertex(h,k)
.
therefore complete the square, x^2 +2x + 1,,,,,(2/2)^2=1
.
add +1 to both sides of eqn
.
y+6+1 = x^2 +2x +1
.
(y+7) = (x+1)^2,,,,,,,,therefore vertex at (-1,-7)
.
positive coefficient of x^2 ,,,therefore opening up,,,y parabola
.
vertex is at a min,,,as pointing up.
.
line of symmetry is at ,,,,x=-1
.
Just check a few points,,(0,-6) and (-2,-6),,,,ok